Session: Oct 2010 (session 3)
Wavelength: 1.3 cm
To give feedback on this experiment, click
Onsala (Michael Lindqvist at Mon Nov 01 05:29:11 2010) Success:
No known problems (I guess this should be EA044).
Yebes (Paco Colomer at Mon Nov 01 08:32:11 2010) Success:
No known problems. Feedback refers to experiment EA044.
(Andrea Orlati at Tue Nov 02 07:28:50 2010) Success:
No known problems
Effelsberg (Uwe Bach at Tue Nov 02 08:50:38 2010) Minor failures:
Missed first 10 minutes because of antenna control problems.
Metsahovi (Minttu Uunila at Thu Nov 04 09:15:30 2010) Minor failures:
BBC01 is suspected to be broken although reports 'lock'.
Metsahovi (Minttu Uunila at Thu Nov 04 09:41:50 2010) Minor failures:
BBC02 hopefully locked to mid-range ~750MHz line setting.